OpenStudy (anonymous):

please solve !! -4l8-xlle-20

OpenStudy (anonymous):

please solve !! -4l8-xl le-20

OpenStudy (anonymous):

\[-4\left| 8-x \right|\le-20\]

OpenStudy (s3a):

2 secs

OpenStudy (s3a):

is it the absolute value that's confusing you? the inequality? or both?

myininaya (myininaya):

i would divide both sides by -4 first (doing this you have to flip the inequality sign)

OpenStudy (anonymous):

i need to solve it

OpenStudy (s3a):

just to be clear, you flip the inequality sign whenever multiplying by a negative number and division is considered multiplication of a reciprocal. the reciprocal of a is 1/a for example

OpenStudy (s3a):

is it the absolute value that's confusing you? the inequality? or both?

myininaya (myininaya):

\[|f(x)|>a => f(x)<-a \text{ or } f(x)>a\]

OpenStudy (anonymous):

hmmm ok so the result would be?? x=13???

myininaya (myininaya):

\[|8-x| \ge 5\]

myininaya (myininaya):

\[8-x \le -5 \text{ or } 8-x \ge 5\]